OpenStudy (anonymous):

Find the value of 3 over the square root of 3..

OpenStudy (anonymous):

\[\frac{a}{\sqrt{a}}=\frac{a}{\sqrt{a}}\times \frac{\sqrt{a}}{\sqrt{a}}=\frac{a\sqrt{a}}{a}=\sqrt{a}\]

OpenStudy (anonymous):

you can check that it is right, i.e. that \[\frac{a}{\sqrt{a}}=\sqrt{a}\] because \(\sqrt{a}^2=a\) and also \[\left(\frac{a}{\sqrt{a}}\right)^2=\frac{a^2}{a}=a\]

OpenStudy (saifoo.khan):

@satellite73 is on fire!

OpenStudy (hba):

\[3=\sqrt{3 \times 3}\]